How do you find the value of x so that l is parallel to m
3x+ 20 is on the top line on the left side and 2x+40 is on the second line on the left. Its a transversal line.

/
-----/-----
/
---/-------
/

Answer by gonzo(654)

i'm guessing, but i think this is what you want.
-----
you have 2 lines that are parallel.
you call them l and m.
i'll call them
l = AB (goes from point A on the left to point B on the right).
m = CD (goes from point C on the left to point D on the right).
line l is above line m.
-----
you have a third line that cuts across l and m (it is a transversal.
i'll call that line k equal to GH (goes from point G on the bottom to point H on the top).
based on the answer to this problem, if correct, the line slants slightly to the left going from bottom (point G) to top (point H).
-----
i'll call the intersection of line k with line l the point E.
i'll call the intersection of line k with line m the point F.
-----
i believe you are referring to angle measures when you say (3x+20) and you say (2x+40).
-----
if i understand you correctly, then angle AEH equals 3x+20, and angle CFH = 2x+40
-----
that would make these angles corresponding angles of parallel lines and so they must be equal to each other.
-----
if that's true, then
3x+20 = 2x+40
subtract 2x from both sides:
3x-2x+20 = 40
subtract 20 from both sides:
3x-2x = 40-20
simplify:
x = 20
-----
the value of x = 20 will make both these angles equal.
3(20) + 20 = 60 + 20 = 80
2(20) + 40 = 40 + 40 = 80
-----
since those angles are 80 degrees, the line k is slanting slightly to the left going from bottom to top (point G to point H).
-----